# Language Model Evolutionary Algorithms for Recommender Systems: Benchmarks and Algorithm Comparisons

### v1

@@ IV-B3 LLM-IBEA
0:  The nondominated recommendation prompt set $\mathcal{P}$;
1:  Generate the initial population $\mathcal{P}=\{\textbf{p}_{1},\ldots,\textbf{p}_{N}\}$ by using the initial operations introduced in Section IV-A1;
2:  Evaluate the initial population by using the objective functions;
3:  $t\leftarrow 0$
4:  while $t<t_{max}$ do
5:     Generate an offspring population of recommendation prompts $\mathcal{Q}=\{\textbf{q}_{1},\ldots,\textbf{q}_{N}\}$ according to the operations introduced in Section IV-A2;
6:     Evaluate the offspring population by using the objective functions;
7:     $\mathcal{P}\leftarrow\mathcal{P}\cup\mathcal{Q}$;
8:     Select $N$ recommendation prompts from $\mathcal{P}$ based on the $I_{(\epsilon)^{+}}$ indictor based environment selection, and forming a new generation of population;
9:     $t\leftarrow t+1$;
10:  end while
11:  return  A population of nondominated recommendation prompts $\mathcal{P}$.
@@
Problems Training Set Validation set
@@
RSBench-1 0.6606$\pm$0.0145 0.6652$\pm$0.0113 0.7088$\pm$0.0045 0.4978$\pm$0.0350 0.4877$\pm$0.0188 0.5186$\pm$0.0106
RSBench-2 0.5901$\pm$0.0333 0.6219$\pm$0.0270 0.6694$\pm$0.0386 0.4155$\pm$0.0198 0.3951$\pm$0.0254 0.4413$\pm$0.0570
RSBench-3 0.7637$\pm$0.0691 0.7578$\pm$0.0134 0.7448$\pm$0.0235 0.5308$\pm$0.0382 0.5122$\pm$0.0225 0.4302$\pm$0.0293
RSBench-4 0.4169$\pm$0.0284 0.5586$\pm$0.0411 0.5137$\pm$0.0413 0.3302$\pm$0.0592 0.3877$\pm$0.0217 0.3551$\pm$0.0338
RSBench-5 0.4143$\pm$0.0263 0.3615$\pm$0.0160 0.4171$\pm$0.0220 0.2765$\pm$0.0406 0.2433$\pm$0.0230 0.2400$\pm$0.0271
RSBench-6 0.2680$\pm$0.0273 0.2451$\pm$0.0211 0.2842$\pm$0.0231 0.1426$\pm$0.0271 0.1308$\pm$0.0165 0.1523$\pm$0.0244
RSBench-7 0.3511$\pm$0.0113 0.3914$\pm$0.0277 0.4116$\pm$0.0281 0.2254$\pm$0.0208 0.2311$\pm$0.0235 0.2142$\pm$0.0270
RSBench-8 0.2834$\pm$0.0238 0.2509$\pm$0.0079 0.2676$\pm$0.0250 0.1510$\pm$0.0239 0.1341$\pm$0.0151 0.1597$\pm$0.0229
RSBench-9 0.2580$\pm$0.0291 0.2724$\pm$0.0207 0.2652$\pm$0.0318 0.1186$\pm$0.0232 0.1225$\pm$0.0204 0.1250$\pm$0.0227
Figure 6: Pareto front approximation results provided by LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA on RSBench-1, RSBench-5, and RSBench-9 with the randomly sampled batch in training set. (a) LLM-NSGA-II results on RSBench-1. (b) LLM-MOEA/D results on RSBench-1. (c) LLM-IBEA results on RSBench-1. (d) LLM-NSGA-II results on RSBench-5. (e) LLM-MOEA/D results on RSBench-5. (f) LLM-IBEA results on RSBench-5. (g) LLM-NSGA-II results on RSBench-9. (h) LLM-MOEA/D results on RSBench-9. (i) LLM-IBEA results on RSBench-9.
@@ V-A Parameter Settings and Implementation Details
In our experiment studies, GPT-3.5 [7] is employed in both the RSBench and the LLM-based EAs. For the algorithms LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA, we set the maximum number of iterations to 20 and the population size to 10. In LLM-MOEA/D, 10 weight vectors are generated using the Riesz s-energy method [72], with a neighborhood size of 5. The environmental selection mechanisms for LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA adhere to those specified in the original papers [33, 36, 38]. In each iteration, a batch of 10 samples is randomly selected from the training set to evaluate objective function values. For each sample, 20 items are saved in the candidate set; if the target item ranks within the top 10 after re-ranking by the RS, the recommendation is deemed correct. The optimization process is repeated five times, each with a different random seed (i.e., 0, 10, 42, 625, and 2023) to generate diverse candidate sets. For the validation set evaluation, we randomly select 50 samples to estimate the objective functions, maintaining all other settings consistent with those used for the training set.
@@ V-B Hypervolume Measure
We evaluated the algorithm quality using the hypervolume (HV) metric, which measures the size of the objective space region dominated by a set of solutions [73]. Since the objective values in the RSBench problems are maximization problems and confined to the region $[0,1]$, we use the point $(0,0)$ as the reference for two-objective problems (i.e., RSBench-1 to RSBench-6) and $(0,0,0)$ as the reference for three-objective problems (i.e., RSBench-7 to RSBench-9).
V-C Results of LLM-Based EAs on RSBench
@@
Figure 7: Pareto front approximation results provided by LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA on RSBench-1, RSBench-5, and RSBench-9 with the randomly sampled batch in validation set. (a) LLM-NSGA-II results on RSBench-1. (b) LLM-MOEA/D results on RSBench-1. (c) LLM-IBEA results on RSBench-1. (d) LLM-NSGA-II results on RSBench-5. (e) LLM-MOEA/D results on RSBench-5. (f) LLM-IBEA results on RSBench-5. (g) LLM-NSGA-II results on RSBench-9. (h) LLM-MOEA/D results on RSBench-9. (i) LLM-IBEA results on RSBench-9.
@@
In this subsection, we assess the performance of LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA on RSBench problems. Table III presents the HV results achieved by these algorithms across multiple candidate sets generated with different random seeds. The results reveal that LLM-IBEA exhibits the best overall performance, achieving the highest HV values on five RSBench problems (RSBench-1, RSBench-2, and RSBench-5 to RSBench-7) with the training sets and on five problems (RSBench-1, RSBench-2, RSBench-5, RSBench-8, and RSBench-9) with the validation sets. LLM-NSGA-II and LLM-MOEA/D show comparable performance, with each achieving the highest HV results on two RSBench problems with both the training and validation sets. Specifically, LLM-NSGA-II performs best on RSBench-3 and RSBench-8 with the training set, and on RSBench-3 and RSBench-5 with the validation set. LLM-MOEA/D achieves top performance on RSBench-4 and RSBench-9 with the training set, and on RSBench-4 and RSBench-7 with the validation set.
Additionally, Table IV presents the time each algorithm took to solve each problem. The results indicate that optimizing RSBench problems with LLM-based EAs is notably time-intensive, with each optimization run taking between approximately 2.5 and 8 hours. This high time consumption is primarily due to frequent interactions with the LLM for objective function evaluations in the RSBench problems. Both the inherent latency of LLM inference and the overhead from repeated communication contribute to the significant time costs associated with these evaluations.
V-D Effectiveness of the LLM-Based Initialization
@@
V-E Effectiveness of the LLM-Based Offspring Generation
@@
As described in Section IV-A2, the LLM-based EAs in this paper generate offspring recommendation prompts primarily through LLM-driven crossover and mutation operators. An important question to investigate is whether these LLM-driven crossover and mutation operators truly enhance algorithm performance on RSBench problems. To address this, we examine the performance of LLM-IBEA by comparing it with a variant named LLM-IBEA-wo-C&M, which omits the LLM-driven crossover and mutation operations. In LLM-IBEA-wo-C&M, offspring are generated directly based on all recommendation prompts stored in the population. The generation process is guided by the prompt: ‘Please generate a new prompt based on the following set: {population}. Provide a final prompt bracketed with <<<START>>> and <<<END>>>.’ We then evaluate the performance of both algorithms on RSBench-1 and RSBench-4.
@@ VII Conclusion
To rigorously assess the performance and applicability of LLM-based EAs, real-world benchmark problems are crucial. This paper introduces RSBench, a benchmark problem set designed to evaluate LLM-based EAs in the context of recommendation prompt optimization, with a focus on session-based RS with the LLM. RSBench aims to identify a set of Pareto optimal prompts that guide the recommendation process, delivering recommendations that are accurate, diverse, and fair. Moreover, by integrating an LLM-based initialization strategy along with LLM-driven offspring generation strategies, we developed three LLM-based EAs—LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA—within established multiobjective evolutionary frameworks. Performance evaluations on RSBench reveal that LLM-IBEA achieves the best results across the benchmark problems. Furthermore, our analysis highlights the effectiveness of the LLM-based initialization and LLM-driven offspring generation in enhancing algorithmic performance. We also discuss future research directions that RSBench opens for LLM-based recommendation systems and EAs. We believe RSBench will support and drive future advancements in LLM-based EAs and RSs. Future work will focus on exploring more effective LLM-based EAs and recommendation methodologies, further advancing the application of evolutionary computing in the domain of recommendation systems.
Acknowledgment
The authors acknowledge the use of GPT-4o to identify improvements in the writing, prompts used with “Please refine and optimize the description of the following, and make it be suitable for a research paper”.

### v3

@@ IV-B3 LLM-IBEA
0: The population size NN; objective functions $\textbf{F}(\textbf{p})$;
0: The nondominated recommendation prompt set $\mathcal{P}$;
1: Generate the initial population $\mathcal{P}=\{\textbf{p}_{1},\ldots,\textbf{p}_{N}\}$ by using the initial operations introduced in Section IV-A1, and evaluate $\mathcal{P}$ by using the objective functions;
@@
3:while $t<t_{max}$ do
4:  Generate an offspring population of recommendation prompts $\mathcal{Q}=\{\textbf{q}_{1},\ldots,\textbf{q}_{N}\}$ according to the operations introduced in Section IV-A2;
5:  Utilizing the environmental selection of IBEA to select recommendation prompts from $\mathcal{P}\cup\mathcal{Q}$, thus forming the new population $\mathcal{P}$;
6:  $t\leftarrow t+1$;
7:end while
8:return A population of nondominated recommendation prompts $\mathcal{P}$.
@@
Problems Training Sets Validation Sets
@@
[…]
Figure 6: Pareto front approximation results provided by LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA on RSBench-1, RSBench-5, and RSBench-9 with the randomly sampled batch in training set. (a) Results on RSBench-1. (b) Results on RSBench-5. (c) Results on RSBench-9.
@@ V-A Parameter Settings and Implementation Details
In our experiment studies, GPT-3.5 (version gpt-3.5-turbo[7] is employed in both the RSBench and the LLM-based EAs, and the temperature of the LLM is set to 0.7. For the algorithms LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA, we set the maximum number of iterations to 20 and the population size to 10. In LLM-MOEA/D, 10 weight vectors are generated using the Riesz s-energy method [75], with a neighborhood size of 3. (The influence of the neighborhood size is investigated in Section V-G.) The environmental selection mechanisms for LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA adhere to those specified in the original papers [36, 39, 41]. To evaluate the performance of the algorithms, we partitioned each dataset into training and validation sets. Specifically, we randomly split the data in an 8:2 ratio, with 80% allocated to the training set and the remaining 20% to the validation set. In each iteration, a batch of 10 samples is randomly selected from the training set to evaluate objective function values. For each sample, 20 items are saved in the candidate set; if the target item ranks within the top 10 after re-ranking by the RS, the recommendation is deemed correct. The optimization process is repeated five times, each with a different random seed (i.e., 0, 10, 42, 625, and 2023) to generate diverse candidate sets. For the validation set evaluation, we randomly select 200 samples to estimate the objective functions, maintaining all other settings consistent with those used for the training set. Both the RSBench and the LLM-based EAs were implemented in a Windows 10 Professional environment, using an Intel(R) Xeon(R) CPU ES-1650 with a clock speed of 3.6 GHz and 64 GB of RAM. The implementation was carried out using Python version 3.10.2.
@@ V-B Hypervolume Measure
We evaluated the algorithm quality using the hypervolume (HV) metric [76], which measures the size of the objective space region dominated by a set of solutions. Since the objective values in the RSBench problems are maximization problems and confined to the region $[0,1]$, we use the point $(0,0)$ as the reference for two-objective problems (i.e., RSBench-1 to RSBench-6) and $(0,0,0)$ as the reference for three-objective problems (i.e., RSBench-7 to RSBench-9).
V-C Results and Insights of LLM-Based EAs on RSBench
@@
Figure 7: Pareto front approximation results provided by LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA on RSBench-1, RSBench-5, and RSBench-9 with the randomly sampled batch in validation set. (a) Results on RSBench-1. (b) Results on RSBench-5. (c) Results on RSBench-9.
In this subsection, we evaluate the performance of LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA on RSBench problems. Table III presents the HV results obtained by these algorithms across multiple candidate sets generated with different random seeds. The results indicate that LLM-IBEA achieves the best overall performance, attaining the highest HV values on six RSBench problems (RSBench-1, RSBench-2, RSBench-5 to RSBench-7, and RSBench-9) with the training sets and on four problems (RSBench-1, RSBench-2, RSBench-5, and RSBench-6) with the validation sets. LLM-NSGA-II and LLM-MOEA/D exhibit comparable performance. LLM-NSGA-II attains the highest HV value on one RSBench problem in the training set and two in the validation set, specifically excelling on RSBench-8 (training set) and RSBench-3 and RSBench-8 (validation set). LLM-MOEA/D achieves the highest HV values on two RSBench problems in the training set (RSBench-3 and RSBench-4) and three in the validation set (RSBench-4, RSBench-7, and RSBench-9).
In addition to the above experimental results, several key insights have been gained to better inform algorithm design in this area:
  • Decomposition-based methods exhibit relatively unstable performance compared to the other two approaches. A possible explanation is that crossover and mutation operations are conducted within subpopulations that lack sufficient diversity. As a result, the LLM encounters difficulties in generating novel prompts through the recombination of candidate prompts within these limited diverse subpopulations, which restricts the algorithm’s exploration capability. This reduced exploratory ability may cause the prompts to partially converge to suboptimal solutions in certain runs, leading to instability in the optimization outcomes. To examine this, we measure the diversity of prompts in the final population using the BERTScore. Specifically, we compute the BERTScore for each pair of recommender prompts in the final population and use the average score as a diversity metric. The results indicate that LLM-NSGA-II and LLM-IBEA achieve average scores of 0.1005 and 0.1025, respectively, whereas LLM-MOEA/D yields a larger score of 0.1036, reflecting reduced population diversity.
  • Moreover, from the HV results obtained by the three algorithms, we found that the benchmark based on the Bundle dataset proves more difficult to optimize. This may be because RSs relying solely on prompts and LLMs in a zero-shot setting struggle to handle complex data involving diverse items within a bundle.
  • Extending from two-objective to three-objective optimization problems poses additional challenges for all algorithms. This difficulty likely arises from the increased complexity of balancing trade-offs among accuracy, diversity, and fairness simultaneously. This may be obversed from Fig. 6 and Fig. 7. The obtained solutions exhibit an approximately degenerate Pareto front in the target space, making it difficult to generate recommendation prompts that explore regions with low fairness and diversity. This phenomenon is also a characteristic of RSBench with three objectives. In general, if a recommendation result achieves high fairness, less popular items are more likely to appear in the recommendation list. When these rare items are recommended alongside popular ones, the results typically also exhibit high diversity. This positive correlation arises because fairness promotes exposure of underrepresented items, while diversity is naturally enhanced when both frequent and infrequent items are included in the same recommendation list.
Moreover, from a recommendation perspective, a unique advantage of multiobjective evolutionary algorithms lies in their ability to generate a diverse set of prompts that capture trade-offs across different objectives. This diversity enables users to flexibly select the prompts that best align with their preferences, offering greater adaptability compared to non-EA-based prompt optimization strategies, which typically provide fewer options. We also compared two representative prompts obtained from LLM-IBEA to examine their impact on recommendation outcomes. We first give the two representative prompts:
  • Prompt A (higher diversity, lower accuracy): Identifying unique combinations of items present in the session is crucial. Analyze the patterns of sequential user actions within the session and identify any recurring sequences. We analyze the items within each combination to deduce the user’s specific interaction intent for that combination. Next, we determine the most appropriate user intent from the inferred intents that align with the user’s current preferences. Predict the user’s next intended action based on the patterns observed in the sequential data. Utilizing the selected intent, we reevaluate the 20 items in the candidate set for potential user interactions and present the revised ranking results including the item indices while maintaining the order of all items in the candidate set exclusively. Determine the most prominent user intent based on the analysis of patterns and interactions. Determine the likelihood of successful completion for each predicted action. Reorder the 20 items in the candidate set according to the likelihood of alignment with the identified intent, providing the revised ranking results with item indices. Prioritize the predicted actions based on the likelihood of successful completion and present the top 3 recommendations to the user.
  • Prompt B (higher accuracy, lower diversity): Identifying patterns in the user’s session interactions to reveal any recurring themes or preferences is crucial. We need to analyze the items within each combination to deduce the user’s specific interaction intent for that combination. Determine the most relevant intent from the inferred ones that align with the user’s current preferences. Next, we analyze the relationship between the identified patterns and the items accessed to understand the user’s intent behind each interaction. Utilizing the selected intent, we reorder the 20 items in the candidate set based on the likelihood of alignment with the identified intent, providing the revised ranking results with item indices. Ensure that the order of all items in the candidate set is maintained, and the items for ranking are exclusively from the candidate set.
Prompt A emphasizes repeated inference of user intent and multiple re-ranking operations within the candidate set. This design introduces broader coverage of potential user preferences and consequently increases the diversity of the recommendations. However, the repeated reasoning and ranking steps also add instability and noise, which can reduce accuracy. In contrast, Prompt B adopts a more streamlined structure. It focuses on identifying a single most relevant intent and directly reorders the candidate set based on the likelihood of alignment with this intent. By limiting the process to one intent extraction and strictly constraining the output to the candidate set, Prompt B yields more accurate recommendations. Nevertheless, this design reduces the exploratory capacity of the model, resulting in lower diversity in the final outcomes.
V-D Time and Token Consumptions on RSBench Problems
@@
TABLE V: Consumed tokens of LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA on RSBench Problems, averaged over 5 Independent Runs with Diverse Generated Candidate Sets.
Problems LLM-NSGA-II LLM-MOEA/D LLM-IBEA
Average Consumed Tokens
RSBench-1 2831028 2249745 2593793
RSBench-2 3739494 3312846 3291100
RSBench-3 5017110 4994992 5106746
RSBench-4 2909850 2236588 3442009
RSBench-5 3901471 2916929 3574600
RSBench-6 6120710 4377109 5981138
RSBench-7 2814915 2435063 2715296
RSBench-8 3303001 2683367 2989405
RSBench-9 5279394 5148334 5192066
In the optimization process of RSBench, frequent calls to the LLM’s API are required for both optimization and evaluation, leading to significant time and token consumption. In this subsection, we further analyze the time and token consumption during the optimization process. Table IV presents the time each algorithm took to solve each problem. The results indicate that optimizing RSBench problems with LLM-based EAs is notably time-intensive, with each optimization run taking between approximately 2.5 and 8 hours. This high time consumption is primarily due to frequent interactions with the LLM for objective function evaluations in the RSBench problems. Both the inherent latency of LLM inference and the overhead from repeated communication contribute to the significant time costs associated with these evaluations. Moreover, since the algorithms rely on calls to the GPT-3.5 API, the call time is influenced by the API provider. Therefore, it is more reasonable to compare the average number of tokens consumed across different algorithms. To facilitate this comparison, we also present the token consumption of the algorithms in Table V.
V-E Effectiveness of the LLM-Based Initialization
@@
V-F Effectiveness of the LLM-Based Offspring Generation
@@
As described in Section IV-A2, the LLM-based EAs in this paper generate offspring recommendation prompts primarily through LLM-driven crossover and mutation operators. An important question to investigate is whether these LLM-driven crossover and mutation operators truly enhance algorithm performance on RSBench problems. To address this, we examine the performance of LLM-IBEA by comparing it with a variant named LLM-IBEA-wo-C&M, which omits the LLM-driven crossover and mutation operations. In LLM-IBEA-wo-C&M, offspring are generated directly based on all recommendation prompts stored in the population. The generation process is guided by the prompt: ‘Please generate a new prompt based on the following set: {population}. Provide a final prompt bracketed with <<START>> and <<END>>.’ We then evaluate the performance of both algorithms on RSBench-1 and RSBench-4.
@@
V-G Influence of the Neighborhood Size of LLM-MOEA/D
TABLE VI: Results of LLM-MOEA/D with the Neighborhood Size of 3, 5, and 7.
Problems LLM-MOEA/D with Different Neighborhood Sizes
0.3 0.5 0.7
[…]
In this subsection, we analyze the impact of neighborhood size on the performance of LLM-MOEA/D. Specifically, we evaluate the algorithm with neighborhood sizes of 3, 5, and 7 on RSBench-1 and RSBench-2. The results indicate that, for both RSBench-1 and RSBench-2, a neighborhood size of 3 yields the best performance. Therefore, we adopt this setting for the remaining experiments.
@@ VII Conclusion
To rigorously assess the performance and applicability of LLM-based EAs, real-world benchmark problems are essential. This paper primarily discusses the following three key points:
  1. 1.
    We introduce a benchmark problem set, RSBench, designed to evaluate LLM-based EAs in the context of recommendation prompt optimization, with a focus on session-based RS utilizing LLMs. RSBench aims to identify a set of Pareto-optimal prompts that guide the recommendation process, delivering recommendations that are accurate, diverse, and fair.
  2. 2.
    We integrate an LLM-based initialization strategy with LLM-driven offspring generation strategies to develop three LLM-based EAs—LLM-NSGA-II, LLM-MOEA/D, and LLM-IBEA—within established multiobjective evolutionary frameworks. Performance evaluations on RSBench demonstrate that LLM-IBEA achieves the best results across the benchmark problems.
  3. 3.
    We conduct experimental investigations to evaluate the effectiveness of LLM-based initialization and LLM-driven offspring generation in enhancing the overall performance of the algorithms.
In addition to these three points, we also discuss potential future research directions that RSBench opens for LLM-based recommendation systems and EAs. We believe RSBench will significantly contribute to advancing LLM-based EAs and recommendation systems. Further work will focus on developing more effective objective estimation methods to improve the performance of LLM-based EAs in RSs, designing more efficient EA frameworks to tackle complex optimization problems in RSs, and exploring the potential of LLM-based EAs in enabling more powerful recommended capabilities in RSs.